OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@goformit100 help plz

OpenStudy (goformit100):

@Aperogalics

OpenStudy (anonymous):

ok if we rearrange it it will become comparing it with y=mx+c we get slope m=-3 so a line perpendicular to this line must have slope 'M' such that mM=-1 M=-1/-3=1/3 so y-y1=m(x-x1) is the equation of line if m=slope and (x1,y1)=point given